1. Suppose Mary intends to sell two software products X & Y for the next convention &


budgets the following.


X Y Total


Units Sold. 60 40 100


Revenues, $200 $100 per unit $12,000 $ 4,000 $16,000


Variable Costs, $120 $70 per unit 7,200 2,800 10,000


Unit Contribution Margin, $80 $ 30 per unit $ 4,800 $ 1200 $ 6,000


Fixed Costs 4,500


Operating Income $ 1,500


Required: What is the BEP (in units & in Birr


Required: Answer the following


1. What is the sales mix of videos and equipment sets


2. Compute weighted average contribution margin


3. Compute the break-even quantity of each product.


4. What is weighted average contribution margin ratio


5. What is the overall break-even sales revenue


1
Expert's answer
2022-03-15T04:30:03-0400
  1. $4500
  2. (80+30+4800+1200+600)/5 = 1342
  3. 70-30= 40USD
  4. 80+30+4800+1200+600)/5 = 1342
  5. 4500-1500 = 3000.